a boy sitting on a swing at a maximum height of 5m above the ground.when the swing passes through the mean position which is 2m above the ground its velocity is?
1
Expert's answer
2012-10-05T07:55:38-0400
The law of conservation of energy states that the total amount of
energy in
an isolated system remains constant over time.
At a maximum height:
K0 =
0
P0 = m * g * H
At a lowest point:
K1 = m * V^2 / 2
P1 = m * g *
h

K0 + P0 = K1 + P1, so:
m * g * H = m * V^2 / 2 + m * g * h
g * H
= V^2 / 2 + g * h
V^2 / 2 = g * H - g * h
V = sqrt(2 * g * (H - h))
V =
sqrt(2 * 9.81 * (5 - 2)) = 7.67 m/s